7. Use the Logarithmic laws to simplify y first and then find the derivative y. 5 x +3 (a) y = log1 (3+5)4 (b) y = log (1+)8. Use logarithmic differentiation to find y if y = 3+ (x - 2) x41x + 5 (b) y = a. 12.0 (C) y = (1+ (Use the result of Q7(b)
Answer #1

x+3 ox, y = bang ( 59J348 ) - bag (et 53€ or, y = { lng 5* + heng Jass} - 46g (at 5) oks g = x log 5 + 1 lug (x+3) - F log (xACCUEIL X +1 Y I yappa stor + 2 baybay-fe of t 2 lapas | org y = 2 ad + 2 lag (8+1) - 2 = 2 loga , y = ( -2) + (2 bag Get30 = * ок, 4 = 4 (***) ok, log y = ex lega Differentiating bath sédes wird a, we get 58 = ** *+ leax e**. 2х | Ф, та = е** (on, y= (1++) 2* * * + 2 bag(et) 2 -2 logas Ok, yo = (1+2) * { 2x + log(x+1) 7 -2- log 2 ²} Rt!